Find the compound amount and compound interest of principal Rs 10000 for 2 years at the rate of 6% p.a without using formula.

Grade:10

1 Answers

Arun
4 years ago
Dear student
 
for 1st year 10000* 6/100 = 600 ₹
 
now for second year
 
10600 * 6/100 = 636 ₹
 
hence compound interest = 600 + 636 = 1236 ₹
compound money = 1000 + 1236 = 2236 ₹